# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

### 18 Mar 2026 v2

@@ 5.1.3 Jackknife Consistency
We establish (a) and (b) separately, then combine via the $L^{2}$ criterion $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$.
(a) Expectation: $\mathbb{E}[C_{n}]\to c_{R}$.
Step (a1): Reduce to $b_{i}$ terms. From Lemma 5.1, $\delta_{i}=a_{i}+b_{i}$ with $|a_{i}|\leq M/n$ deterministically and $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$. Since $|\delta_{i}^{2}-b_{i}^{2}|=|a_{i}||2b_{i}+a_{i}|\leq(M/n)(2|b_{i}|+M/n)$,
$\bigl|\mathbb{E}[C_{n}]-(n-1)\textstyle\sum_{i}\mathbb{E}[b_{i}^{2}]\bigr|\;\leq\;2Mn\cdot O(r_{n}/n^{1/2})\cdot(M/n)+O(n^{-1})=O(n^{-1}),$ (12)
so it suffices to show $(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]\to c_{R}$.
Step (a2): Apply Assumption 5.1. By (14), $(n-1)\sum_{i}b_{i}^{2}\xrightarrow{P}c_{R}$ already. Taking expectations, justified by uniform integrability condition (iii) of Assumption 5.1 and the bound $|b_{i}|\leq 2M/n$ on $\mathcal{E}_{n}$ (which holds with probability tending to 1): $\mathbb{E}[C_{n}]\to c_{R}$. $\square_{\text{(a)}}$
@@
The LOO remainder perturbations $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$ admit a uniform linearization: there exist mean-zero random variables $\{\psi_{n,i}\}_{i=1}^{n}$ such that
$\max_{1\leq i\leq n}\bigl|nb_{i}+\psi_{n,i}\bigr|\;\xrightarrow{P}\;0,$ (13)
and the sequence $\{\psi_{n,i}\}$ satisfies the following moment conditions:
  1. i.
    (Negligible mean): $n^{-1}\sum_{i=1}^{n}\psi_{n,i}=o_{P}(1)$.
  2. ii.
    (Variance identification): $n^{-1}\sum_{i=1}^{n}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$, where $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})>0$.
  3. iii.
    (Uniform integrability): $\sup_{n}\mathbb{E}\bigl[\max_{i}\psi_{n,i}^{2}\bigr]<\infty$.
Remark 5.1 (Scope and verification of Assumption 5.1).
Assumption 5.1 targets exactly what Proposition 5.1 needs: the identification of the leading second-moment contribution $\mathbb{E}[C_{n}]\to c_{R}$. It assumes the LOO linearization directly, without deriving it from a smoothness condition on the nuisance maps.
A sufficient condition for (13) is Hadamard differentiability of the nuisance maps $P\mapsto\hat{\eta}_{k}(P)$, $k=1,2$, at $P_{0}$ tangentially to the set of signed measures. Under this condition, the functional delta method gives $nb_{i}=-\psi(O_{i})+o_{P}(1)$ uniformly in $i$, where $\psi\in L^{2}_{0}(P_{0})$ is the influence function of $P\mapsto R_{\mathrm{rem}}(P)$ (van der Vaart, 1998). Setting $\psi_{n,i}=\psi(O_{i})$ then satisfies conditions (i)–(iii) by the LLN and $\mathrm{Var}(\psi)=c_{R}$. This holds for regular plug-in estimators and for cross-fitted Super Learner with a Donsker-class library, but is not automatic for arbitrary data-adaptive procedures.
Under additional algorithmic-stability or bounded-differences conditions on the nuisance learners, one may alternatively verify the concentration step $C_{n}-\mathbb{E}[C_{n}]\to_{P}0$ directly via Efron-Stein or related arguments, treating concentration as a sufficient-condition tool rather than the backbone of the proof. This is the role assigned to Part (b) of Proposition 5.1 below.
Assumption 5.1 affects only Proposition 5.1(a). It does not affect the main validity argument of Theorem 5.2, which proceeds via the bootstrap route.
Under Assumption 5.1, the identification step is immediate. From (13) and moment conditions (i)–(ii):
$(n-1)\sum_{i=1}^{n}b_{i}^{2}=\frac{1}{n}\sum_{i=1}^{n}(nb_{i})^{2}+O_{P}(n^{-1})=\frac{1}{n}\sum_{i=1}^{n}\psi_{n,i}^{2}+o_{P}(1)\;\xrightarrow{P}\;c_{R},$ (14)
where the first equality uses $nb_{i}=O_{P}(1)$ uniformly from Lemma 5.1, the second uses (13), and the convergence uses condition (ii). Taking expectations (justified by uniform integrability condition (iii)): $\mathbb{E}[C_{n}]\to c_{R}$.
(b) Concentration: $\mathrm{Var}(C_{n})\to 0$.
Reduction to $b_{i}$ terms. Since $|a_{i}|\leq M/n$ deterministically,
$\Bigl|C_{n}-(n-1)\sum_{i}b_{i}^{2}\Bigr|\;\leq\;(n-1)\sum_{i}|2a_{i}b_{i}+a_{i}^{2}|\;\leq\;(n-1)\Bigl[2\frac{M}{n}\sum_{i}|b_{i}|+n\frac{M^{2}}{n^{2}}\Bigr].$
By Cauchy-Schwarz, $\sum_{i}|b_{i}|\leq\sqrt{n\sum_{i}b_{i}^{2}}=O_{P}(n^{1/2}\cdot n^{-1/2})=O_{P}(1)$, so $C_{n}-(n-1)\sum_{i}b_{i}^{2}=O_{P}(n^{-1})$. It suffices to show $\mathrm{Var}\bigl((n-1)\sum_{i}b_{i}^{2}\bigr)\to 0$.
Applying Assumption 5.1. By the LOO linearization (13), $nb_{i}=-\psi_{n,i}+r_{n,i}$ where $\max_{i}|r_{n,i}|\xrightarrow{P}0$. Therefore:
$(n-1)\sum_{i}b_{i}^{2}=\frac{1}{n}\sum_{i}(nb_{i})^{2}+O_{P}(n^{-1})=\frac{1}{n}\sum_{i}\psi_{n,i}^{2}+\frac{1}{n}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})+O_{P}(n^{-1}).$ (15)
The cross term satisfies $\bigl|n^{-1}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})\bigr|\leq 2\max_{i}|r_{n,i}|\cdot n^{-1}\sum_{i}|\psi_{n,i}|+\max_{i}|r_{n,i}|^{2}\xrightarrow{P}0$, using $n^{-1}\sum_{i}|\psi_{n,i}|=O_{P}(1)$ from condition (ii) and Jensen. Hence $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$.
Variance of the leading term. Under the sufficient condition of Remark 5.1 (Hadamard differentiability), $\psi_{n,i}=\psi(O_{i})$ are i.i.d. with $\mathrm{Var}(\psi^{2})\leq\mathbb{E}[\psi^{4}]<\infty$ (condition (iii)), so
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi(O_{i})^{2}\Bigr)=\frac{\mathrm{Var}(\psi(O_{1})^{2})}{n}\to 0.$
In the general triangular-array case ($\psi_{n,i}$ depending on the full sample through shared nuisance fits), condition on the nuisance training data $\mathcal{T}_{n}$. Given $\mathcal{T}_{n}$, the $\psi_{n,i}^{2}$ are deterministic functions of the individual observations $O_{i}$ evaluated at fixed nuisance fits, hence conditionally independent. The conditional variance satisfies:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr)=\frac{1}{n^{2}}\sum_{i}\mathrm{Var}(\psi_{n,i}^{2}\mid\mathcal{T}_{n})\;\leq\;\frac{\max_{i}\mathbb{E}[\psi_{n,i}^{4}\mid\mathcal{T}_{n}]}{n}\;\xrightarrow{P}\;0,$
using condition (iii). Taking total variance via the law of total variance:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\Bigr)\;\leq\;\mathbb{E}\Bigl[\mathrm{Var}\!\Bigl(\cdot\mid\mathcal{T}_{n}\Bigr)\Bigr]+\mathrm{Var}\!\Bigl(\mathbb{E}\Bigl[\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr]\Bigr).$
The first term $\to 0$ as shown above. The second term equals $\mathrm{Var}(n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}])$; by condition (ii), $n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}]\xrightarrow{P}c_{R}$, so this variance also $\to 0$. Therefore $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$, and hence $\mathrm{Var}(C_{n})\to 0$. $\square_{\text{(b)}}$
Conclusion. Steps (a) and (b) give $\mathbb{E}[C_{n}]\to c_{R}$ and $\mathrm{Var}(C_{n})\to 0$, so $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$, establishing $C_{n}\xrightarrow{L^{2}}c_{R}$ and hence $C_{n}\xrightarrow{P}c_{R}$. ∎
Remark 5.2 (Architecture of the Proof and Correction Note).
Both parts of the proof use Assumption 5.1. Part (a) uses it to identify the limit $c_{R}$ via the LOO linearization $nb_{i}\approx-\psi_{n,i}$ and the variance condition $n^{-1}\sum_{i}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$. Part (b) uses it to establish concentration: the direct variance calculation writes $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$ and then shows $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$ via conditional independence of the $\psi_{n,i}^{2}$ given the nuisance training data, together with condition (iii) (uniform integrability).
An earlier draft of this proof used the Efron-Stein inequality for Part (b), claiming concentration holds without Assumption 5.1. That argument is incorrect: in the near-boundary regime, $C_{n}\xrightarrow{P}c_{R}>0$ implies $\mathbb{E}[C_{n}^{2}]\to c_{R}^{2}>0$, so the Efron-Stein bound $\mathrm{Var}(C_{n})\leq\sum_{i}\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}]$ yields only $O(n^{-1})$ from the per-swap sensitivity calculation, which is trivially consistent with $C_{n}=O_{P}(1)$ and provides no useful information. The obstacle is structural: controlling $\mathrm{Var}(C_{n})$ requires controlling pairwise covariances $\mathrm{Cov}(\delta_{i}^{2},\delta_{j}^{2})$ for $i\neq j$, which depend on the LOO correlation structure captured by Assumption 5.1. The corrected proof replaces Efron-Stein with the direct variance calculation above, reducing the proof length while making the assumption dependence explicit.
@@ 5.1.4 Asymptotic Validity
We derive the jackknife result from bootstrap consistency, avoiding the need to establish a triangular-array CLT for the LOO perturbation array.
@@
$\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})}}\;\Rightarrow\;N(0,1).$ (18)
[…]
Step 2: Numerical equivalence. By Definition 5.3, $\widehat{\mathrm{Var}}_{\mathrm{HC}}=\hat{\rho}\cdot\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ where $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$. Therefore $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ identically for every data realization. This is not an approximation: the HC-corrected interval and the jackknife Wald interval are the same interval.
Step 3: Transfer of validity. Since $\widehat{\mathrm{Var}}_{\mathrm{JK}}\equiv\widehat{\mathrm{Var}}_{\mathrm{HC}}$ and $\widehat{\mathrm{Var}}_{\mathrm{HC}}$ is the jackknife-inflated sandwich, it suffices to show that $\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})/\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\xrightarrow{P}1$. In the near-boundary regime, both $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ and $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ are consistent for the total variance $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$: $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ by Theorem 5.3, and $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ by Theorem 5.1(ii) (which now holds without additional assumption, by Proposition 5.1). Therefore $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$, and by Slutsky’s theorem applied to (18):
$T_{n}=\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}}}=\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{Boot}}}}\cdot\sqrt{\frac{\widehat{\mathrm{Var}}_{\mathrm{Boot}}}{\widehat{\mathrm{Var}}_{\mathrm{JK}}}}\;\Rightarrow\;N(0,1)\cdot 1=N(0,1).$
Remark on the strong decay regime. When $c_{R}=0$, both $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ and $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ converge to $\sigma^{2}_{\mathrm{EIF}}/n$, and $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\to 1$ by Theorem 5.1(i) and Theorem 5.3. The argument is identical. ∎
Remark 5.4 (Direct Proof via LOO CLT: An Open Question).
The proof above derives jackknife validity from bootstrap consistency via the numerical equivalence $\widehat{\mathrm{Var}}_{\mathrm{JK}}\equiv\widehat{\mathrm{Var}}_{\mathrm{HC}}$ and the ratio $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\to 1$. A direct proof would instead establish $N(0,1)$ convergence of $T_{n}$ from the LOO expansion of Lemma 5.1 alone, without invoking the bootstrap.
[…]
  1. i.
    $n^{-1/2}\sum_{i}\mathbb{E}[Z_{ni}\mid\mathcal{T}_{n}]=o_{P}(1)$ (conditional mean negligibility);
  2. ii.
    $n^{-1}\sum_{i}\mathrm{Var}(Z_{ni}\mid\mathcal{T}_{n})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}+c_{R}$ (conditional variance stabilization);
  3. iii.
    $n^{-1}\sum_{i\neq j}\mathrm{Cov}(Z_{ni},Z_{nj}\mid\mathcal{T}_{n})\to 0$ (negligible off-diagonal dependence);
  4. iv.
    A conditional Lindeberg condition: $n^{-1}\sum_{i}\mathbb{E}[Z_{ni}^{2}\cdot\mathbf{1}_{|Z_{ni}|>\varepsilon\sqrt{n}}\mid\mathcal{T}_{n}]\to 0$ for all $\varepsilon>0$.
Condition (iii) is the hardest: leave-one-out nuisance fits share n1n-1 training observations, so $Z_{ni}$ and $Z_{nj}$ for $i\neq j$ are correlated through shared training data. This off-diagonal covariance is not eliminated by cross-fitting alone. Establishing (iii) under entropy conditions on the nuisance function class — so that sensitivity to any one observation is $O(n^{-1/2})$ and off-diagonal covariances are therefore $O(n^{-1})$ with a sum that is $O(1/n)\to 0$ — is a technically non-trivial extension that we leave as an open problem.
Whether the direct and bootstrap-derived routes ultimately require the same assumptions is an interesting question, since Assumption 5.2 (Mallows-2 bootstrap consistency) itself requires conditions on the nuisance estimator’s behaviour under resampling that are analogous to the LOO stability condition of Lemma 5.1. The two proof strategies are morally equivalent; the difference is that Assumption 5.2 is a well-studied condition with known sufficient conditions in the bootstrap literature.
The only remaining delicate step in the current proof architecture is Step (a4) of Proposition 5.1, where one identifies the leading behaviour of the LOO remainder perturbation via $nb_{i}=n(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}})=-\psi(O_{i})+o_{P}(1)$ for the influence function $\psi$ of $P\mapsto R_{\mathrm{rem}}(P)$. As stated in Assumption 5.1 and Remark 5.1, this step is not implied by the bilinear form of $R_{\mathrm{rem}}$ alone; it follows under additional smoothness of the nuisance maps $P\mapsto\hat{\eta}_{k}(P)$, conditions that are standard for sufficiently regular plug-in learners but not automatic for arbitrary data-adaptive procedures. Critically, this issue affects only Proposition 5.1(a), where it is used to identify the leading second-moment contribution of the LOO remainder perturbations. It does not affect the main asymptotic normality argument of Theorem 5.2, which proceeds via the bootstrap route and does not require this expansion. The direct LOO-CLT route remains an alternative proof strategy and is not required for any stated result in this paper.
@@ 5.2.2 Bootstrap Consistency
Condition (20) is the Mallows-2 (Wasserstein-2) bootstrap consistency condition. The foundational reference is Bickel and Freedman (1981), who establish this condition for pairs bootstrap applied to smooth functionals of the empirical measure under moment conditions.
[…]
Condition (i) ($\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$) holds throughout this paper since $\mathrm{Var}(T_{n})=\sigma^{2}_{\mathrm{EIF}}+c_{R}+o(1)$ is bounded. Condition (ii) requires verifying Cheng’s $L^{2}$ maximal inequality for the ALE in the near-boundary regime; this is satisfied under entropy conditions on the nuisance function class that are standard for cross-fitted Super Learner with a Donsker-class library.

### 23 Mar 2026 v3

@@ 5.1.3 Jackknife Consistency
We show $\mathbb{E}[(C_{n}-c_{R})^{2}]\to 0$ by establishing (a) $\mathbb{E}[C_{n}]\to c_{R}$ and (b) $\mathrm{Var}(C_{n})\to 0$ separately.
(a) Expectation. From Lemma 5.1, $\delta_{i}=a_{i}+b_{i}$ with $|a_{i}|\leq M/n$ deterministically, so $|\mathbb{E}[C_{n}]-(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]|=O(n^{-1})$. Under Assumption 5.1, the LOO linearization gives $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$ via $\max_{i}|nb_{i}+\psi_{n,i}|\xrightarrow{P}0$ and $\|nb_{i}\|_{L^{2}}=O(1)$. Taking expectations (justified by uniform integrability condition (iii)): $(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]\to c_{R}$ by condition (ii).
@@
The LOO perturbations $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$ admit a uniform linearization $\max_{i}|nb_{i}+\psi_{n,i}|\xrightarrow{P}0$ for mean-zero $\{\psi_{n,i}\}$ satisfying: (i) $n^{-1}\sum_{i}\psi_{n,i}=o_{P}(1)$; (ii) $n^{-1}\sum_{i}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$; (iii) $\sup_{n}n^{-1}\mathbb{E}[\max_{i}\psi_{n,i}^{2}]<\infty$ and $\sup_{n,i}\mathbb{E}[\psi_{n,i}^{4}]<\infty$. Additionally, each $\psi_{n,i}$ is a measurable function of $O_{i}$ alone given the nuisance training data $\mathcal{T}_{n}$, so that $\{\psi_{n,i}\}_{i=1}^{n}$ are conditionally independent given $\mathcal{T}_{n}$.
Remark 5.1 (Scope of Assumption 5.1).
Assumption 5.1 is used in both parts of Proposition 5.1: Part (a) uses the linearization to identify $\mathbb{E}[C_{n}]\to c_{R}$; Part (b) uses the conditional independence structure to establish $\mathrm{Var}(C_{n})\to 0$. It is compatible with Hadamard-differentiable regular plug-in estimators; sufficient conditions for highly adaptive ensemble learners are case-specific and discussed in the Supplementary Material. It does not affect Theorem 5.2, which proceeds via the bootstrap route independently of Assumption 5.1.
(b) Concentration. Since $|a_{i}|\leq M/n$, Cauchy–Schwarz gives $C_{n}-(n-1)\sum_{i}b_{i}^{2}=O_{P}(n^{-1})$, so it suffices to show $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$. Under the conditional array structure of Assumption 5.1, the $\psi_{n,i}^{2}$ are conditionally independent given $\mathcal{T}_{n}$ (each is a function of $O_{i}$ at fixed nuisance fits), so the conditional variance is $\leq\max_{i}\mathbb{E}[\psi_{n,i}^{4}\mid\mathcal{T}_{n}]/n\xrightarrow{P}0$ by the fourth-moment part of condition (iii). The second term of the law of total variance also vanishes since $n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}]\xrightarrow{P}c_{R}$ by condition (ii). ∎
Remark 5.2 (Proof note).
The Efron–Stein inequality is insufficient for Part (b) in the near-boundary regime: $C_{n}\xrightarrow{P}c_{R}>0$ implies $\mathbb{E}[C_{n}^{2}]\to c_{R}^{2}>0$, so the per-swap bound yields only the trivially true $\mathrm{Var}(C_{n})=O(1)$. The direct variance calculation uses the conditional independence structure of Assumption 5.1: given $\mathcal{T}_{n}$, the $\psi_{n,i}^{2}$ are independent, so the conditional variance vanishes at rate $n^{-1}$, and the total variance vanishes by the law of total variance.
@@ 5.1.4 Asymptotic Validity
By Theorem 5.3 and Assumption 5.2, the bootstrap pivot $\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))/\sqrt{n\widehat{\mathrm{Var}}_{\mathrm{Boot}}}\Rightarrow N(0,1)$. The numerical identity $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ (Definition 5.3) and the consistency $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$ (from Theorems 5.1 and 5.3, both estimating $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$) then give $T_{n}=T_{n}^{\rm boot}\cdot\sqrt{\widehat{\mathrm{Var}}_{\mathrm{Boot}}/\widehat{\mathrm{Var}}_{\mathrm{JK}}}\Rightarrow N(0,1)$ by Slutsky’s theorem. The argument is identical when $c_{R}=0$. ∎
Remark 5.4 (Alternative direct proof).
@@ 5.2.2 Bootstrap Consistency
Assumption 5.2 is the Mallows-2 (Wasserstein-2) bootstrap consistency condition studied by Bickel and Freedman (1981) and Cheng (2014). The latter show that $L^{2}$ moment consistency requires an additional maximal inequality beyond distributional consistency; their condition is applicable under entropy bounds on the nuisance function class, covering fixed-weight ensembles and smooth sieves; verification for generic adaptive learners is case-specific. Condition (i) holds since $\mathrm{Var}(T_{n})=\sigma^{2}_{\mathrm{EIF}}+c_{R}+o(1)$ is bounded.